1. Give an example of a pair of enverse operations.____________________.
2. the number of flims (y) is a direct variation of the number of flams (x).
a)write a general equation tht sumarizes thes ralationship.
b)if the consgtan of variation is 15, how many flims are produced by 3 flams?

Answer by stanbon(75887)

1. Give an example of a pair of inverse operations.____________________.
f(x) = 3x + 4
f^-1(x) = (x-4)/3 = (1/3)x - (4/3)
------------------------------------------
2. the number of flims (y) is a direct variation of the number of flams (x).
a)write a general equation tht sumarizes thes ralationship.
y = kx
---------------------
b)if the constant of variation is 15, how many flims are produced by 3 flams?
y = 15x
y = 15*3 = 45 flims are produced by 3 flams.
===================
Cheers,
Stan H.